### v1

# ChauBoxplot and AdaptiveBoxplot: two R packages for boxplot-based outlier detection

Abstract
Tukey’s boxplot is widely used for outlier detection; however, its classic fixed-fence rule tends to flag an excessive number of outliers as the sample size grows. To address this limitation, we introduce two new R packages, ChauBoxplot and AdaptiveBoxplot, which implement more robust methods for outlier detection. We also provide practical guidance, drawn from simulation results, to help practitioners choose suitable boxplot methods and balance interpretability with statistical reliability.
Keywords: Box-and-whisker plot, Chauvenet’s criterion, Chauvenet-type boxplot, Fence coefficient, Outlier detection, Sample size
1 Introduction
The box-and-whisker plot, first introduced by ?, is a fundamental and widely used tool in exploratory data analysis. It provides a concise summary of key distributional features, including the median, quartiles, and range, and visually identifies potential outliers. Because of its non-parametric nature and ease of interpretation, the boxplot is commonly applied in fields such as biomedicine, economics, engineering, psychology, and social sciences, as well as in quality control and educational research. A central function of the boxplot is outlier detection, which is performed by defining fences at fixed distances from the first quartile ($Q_{1}$) and third quartile ($Q_{3}$), and flagging any observations outside these fences as potential outliers. Tukey’s original boxplot sets the lower and upper fences at
${\rm LF}=Q_{1}-k\times\mathrm{IQR}~~~~~{\rm and}~~~~~{\rm UF}=Q_{3}+k\times\mathrm{IQR},$
where $\mathrm{IQR}=Q_{3}-Q_{1}$ is the interquartile range, and $k=1.5$ is the standard fence coefficient. Although this outlier detection rule is simple and widely used, it lacks a rigorous theoretical foundation (?, ?, ?, ?, ?). Specifically, applying a fixed value of $k$ disregards the sample size and distributional characteristics, often leading to an excessive number of false positive outlier flags in larger samples.
To address the limitation imposed by a fixed fence coefficient, ? recently introduced the Chauvenet-type boxplot, which provides exact control of the outside rate per observation. For normally distributed data, by combining Tukey’s boxplot with Chauvenet’s criterion (?), their new fence coefficient is given as
$k_{n}^{\rm Chau}=\frac{\Phi^{-1}(1-0.25/n)}{1.35}-0.5,$ (1)
where $\Phi^{-1}$ denotes the standard normal quantile function and $n$ is the size of the data. Moreover, ? reframed the boxplot-based outlier detection as a multiple testing problem, treating each flagged observation as a statistical test against the majority of the data. Their methods provide a precise control over the family-wise error rate, the per-family error rate, or the false discovery rate. These methodological advances are now available in statistical software, with two new R packages—ChauBoxplot and AdaptiveBoxplot—offering convenient tools for practical implementation. On the other hand, the increased variety of options introduces new challenges, such as selecting suitable methods, interpreting error rates appropriately, and balancing simplicity with statistical rigor.
In this short paper, we present a focused review on the two representative R packages for boxplot-based outlier detection. Section 2 introduces ChauBoxplot, which implements the Chauvenet-type boxplot to provide exact control of the outside rate per observation, outlining its statistical foundation and practical application. Section 3 introduces AdaptiveBoxplot, which applies adaptive multiple testing procedures to improve outlier detection thresholds and accuracy. Section 4 presents a comparative study of the reviewed boxplot methods across a range of sample sizes. Section 5 offers practical guidance and recommendations for selecting boxplot-based outlier detection techniques in modern data analysis.
2 The ChauBoxplot Package
The Chauvenet-type boxplot introduces a novel fence coefficient that, on average, excludes half an observation from normally distributed data regardless of the sample size. Instead of using Tukey’s fixed fence coefficient, it employs the sample-size-adjusted alternative in (1) according to Chauvenet’s criterion. This adjustment integrates seamlessly into the standard boxplot construction, offering enhanced statistical reliability without sacrificing visual intuition or workflow convenience. For further technical details and the theoretical foundation of the Chauvenet-type coefficient, readers may refer to ?.
The ChauBoxplot package provides two main user interfaces for creating the Chauvenet-type boxplot. The chau boxplot()function acts as a direct drop-in replacement for base R’s boxplot(), with the only difference being the default $k=1.5$ replaced by $k_{n}^{\rm Chau}=\Phi^{-1}(1-0.25/n)/1.35-0.5$. For users within the ggplot2 ecosystem, the geom chau boxplot() function can also be implemented similarly to geom boxplot(). It provides a dedicated grammar of graphics layer that supports grouped, faceted, and highly customizable plots, making it well suited for advanced data analysis. Specifically, the new boxplot can be generated using the following syntax:
Ψlibrary(ChauBoxplot)
Ψ# For base R:
Ψchau_boxplot(x)
Ψ# For the ggplot2 ecosystem:
Ψggplot(df, aes(y = x, x = group)) + geom_chau_boxplot()
This compatibility allows researchers to adopt the new boxplot method with minimal changes to their existing code or analysis workflows.
The ChauBoxplot package offers comprehensive documentation, including a reference manual, a README file, and a set of vignettes that demonstrate usage with both synthetic and real-world datasets. These resources guide users through fundamental functions as well as advanced features, such as comparing Chauvenet’s criterion with Tukey’s classic rule and highlighting its advantages for large samples or non-normal data. The package is available on CRAN for all major R platforms and, released under the GPL-3 license, is actively maintained by the original authors to ensure it remains aligned with theoretical developments and evolving user needs. By combining statistically sound, the sample-size-adjusted outlier detection with an intuitive interface and thorough support materials, ChauBoxplot empowers researchers and analysts to conduct modern, reproducible data analysis across a broad range of scientific disciplines.
3 The AdaptiveBoxplot Package
The unified framework by ? extends the methodological advances of ? by systematically reframing the outlier detection as a multiple hypothesis testing problem. In this perspective, Tukey’s classic boxplot corresponds to an unadjusted testing procedure with a fixed per-comparison error rate, while the Chauvenet-type boxplot proposed by ? effectively controls the per-family error rate (PFER). Building on this insight, ? developed a generalized pipeline to construct boxplots that control more advanced error metrics including the family-wise error rate (FWER) and the false discovery rate (FDR). This pipeline proceeds in four steps:
  1. 1.
    Robust Estimation: Parameters of the null distribution are estimated using robust measures, typically the sample quartiles.
  2. 2.
    $\boldsymbol{p}$-value Calculation: Each observation is converted into a $p$-value based on the estimated null distribution.
  3. 3.
    Multiple Testing Adjustment: A desired error rate (e.g., FWER or FDR) and a control target $\alpha$ are first selected. Subsequently, a suitable multiple testing procedure is applied to the set of $p$-values to determine an adjusted significance threshold.
  4. 4.
    Fence Construction: The adjusted threshold is transformed back onto the original data scale to define the boxplot fences.
This approach allows for the creation of fences that are adaptive not only to the sample size but also to the distributional characteristics of the data. Control of the FWER ensures a conservative bound ideal for strict error avoidance, while control of the FDR offers a powerful alternative that scales effectively with data complexity.
The AdaptiveBoxplot package implements this pipeline through two primary functions. The holm boxplot() function utilizes Holm’s step-down procedure (?) to control the FWER. This function is particularly well-suited for applications requiring strict conservatism, where avoiding even a single false positive is a priority. In contrast, the bh boxplot() function applies the Benjamini-Hochberg procedure (?) to control the FDR. This method offers a more powerful approach for large-scale or complex datasets, as it adaptively relaxes the fences when strong evidence of outliers is present, thereby maintaining high sensitivity without succumbing to the inflation of false discoveries typical of fixed-threshold methods.
The package interface is designed for simplicity and consistency with base R graphics. Users can generate error-controlled boxplots for both univariate vectors and grouped data frames using the following syntax:
Ψlibrary(AdaptiveBoxplot)
Ψ
Ψ# Controlling the FWER at level 0.05:
Ψholm_boxplot(data, alpha = 0.05)
Ψ# Controlling the FDR at level 0.05:
Ψbh_boxplot(data, alpha = 0.05)
Apart from the inclusion of the alpha argument, which specifies the target error rate, the usage of both holm boxplot() and bh boxplot() is essentially identical to that of the base function boxplot().
The AdaptiveBoxplot package is freely available on CRAN under the GPL-3 license, supported by a concise reference manual that illustrates the function usage. As an actively maintained project, it ensures that these advanced, multiple-testing-based boxplots remain accessible to the R community. By embedding rigorous statistical inference into a familiar graphical tool, AdaptiveBoxplot facilitates more reproducible and defensible outlier detection. For more details on the unified $p$-value pipeline, readers are referred to ?.
4 Illustrating Example
To illustrate the practical differences among the boxplot-based outlier detection methods, we conducted a simulation study comparing four representative approaches: Tukey’s boxplot, the Chauvenet-type boxplot, the FWER (Holm) boxplot, and the FDR (BH) boxplot. For each setting, we generated a dataset of size $n$, consisting of n3n-3 independent observations from the standard normal distribution and three outliers sampled from $N(5,\,0.5^{2})$. The sample sizes considered were $n=50$, 500500, 50005000 and 50,00050{,}000, allowing us to assess the performance across both small and large datasets. The data generation and analysis were implemented as follows:
Ψlibrary(ChauBoxplot)
Ψlibrary(AdaptiveBoxplot)
Ψset.seed(1000)
Ψn <- c(50, 500, 5000, 50000)
Ψm <- 3
Ψoutliers <- rnorm(m, 5, 0.5)
Ψpar(mfrow = c(4,4))
Ψoldpar = par(mar = c(1, 3.5, 2, 0.5), mgp = c(2, 0.5, 0))
Ψfor (i in 1:length(n)){
Ψ  x <- c(rnorm(n[i]-m), outliers)
Ψ  boxplot(x, main = "Tukey’s boxplot", ylab=paste("n =", n[i]))
Ψ  chau_boxplot(x)
Ψ  holm_boxplot(x, alpha = 0.05)
Ψ  bh_boxplot(x, alpha = 0.05)
Ψ}
Figure 1: Comparison of boxplot-based outlier detection methods for normal data with three contaminated outliers. Each panel displays the results for one method (Tukey, Chauvenet-type, FWER [Holm] with rate 0.05, or FDR [BH] with rate 0.05) at $n=50$, 500500, 50005000 and 50 00050\,000, respectively.
Figure 1 offers a comparative overview of the four boxplot-based outlier detection methods across a range of sample sizes. Each row represents a different sample size, and each column depicts a distinct method, with detected outliers shown as points beyond the whiskers. The three artificially introduced outliers are clearly visible in most panels. As the sample size increases, the classical boxplot with $k=1.5$ tends to misclassify an increasing number of central observations as outliers, especially for n=5,000n=5{,}000 and 50,00050{,}000, where many non-outlying values are incorrectly flagged. In contrast, the Chauvenet-type, FWER (Holm), and FDR (BH) boxplots only flag very few observations in total, demonstrating their robust performance regardless of the sample size.
These results underscore the advantages of sample-size-adjusted and error-controlled boxplots for reliable outlier detection in moderate to large datasets. In contrast, the classical Tukey approach can lead to misleading conclusions when applied to large samples. Our simulation highlights the importance of adaptive strategies, such as those implemented in the ChauBoxplot and AdaptiveBoxplot packages, for principled and reproducible outlier identification.
5 Conclusion and Practical Guidance
In this paper, we reviewed recent advances in boxplot-based outlier detection, focusing on the ChauBoxplot and AdaptiveBoxplot R packages. Our simulation results (see Figure 1) demonstrate that the classic Tukey boxplot increasingly misclassifies regular data points as outliers as the sample size increases. In contrast, the Chauvenet-type boxplot consistently identifies the true outliers in large datasets (e.g., n=50,000n=50,\!000), while minimizing false positives. The FDR-controlled boxplot is also sensitive to outliers, though it may occasionally miss some extreme cases. Meanwhile, the FWER-controlled method is highly conservative and can sometimes overlook genuine outliers, particularly in large samples.
Based on extensive simulation studies (including many results not shown here due to space limitations), we recommend any boxplot method for small datasets, with the FWER-based approach being particularly suitable due to its conservativeness. For moderate-sized datasets (typically 100 to 1,000 observations), the Chauvenet-type and FDR-based boxplots provide the best balance of sensitivity and specificity. In large datasets, the fixed-threshold rules should be avoided; the Chauvenet-type boxplot remains robust across all sample sizes, while FDR-based methods—with a relaxed significance level—offer greater sensitivity when multiple outliers are expected. For all data sizes, robust estimation and careful consideration of the underlying distribution are essential, especially with skewed or non-normal data. Supported by well-documented R packages, these modern approaches enable more accurate, interpretable, and reproducible outlier detection, thereby providing a solid foundation for future advances in statistical graphics and robust data analysis.